40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 438     Entry time: Tue Apr 22 22:19:02 2008
Author: rob 
Type: Metaphysics 
Category: lore 
Subject: jiggling sliders 

In the interests of tacit communication of scientific knowledge, I here reveal a nugget of knowledge which may or may not prove useful to new LIGOites: sometimes when front-end machines are rebooted, the hardware they control can wind up in a state which is not accurately represented by the EPICS values you may see. This can be easily rectified by momentarily changing the EPICS settings in question. For reference, this came up tonight in the context of the whitening gain sliders for the TransMon QPDs.
ELOG V3.1.3-